OpenStudy (anonymous):

What is 30% of 70? i got 21 .... ?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

.3 * 70.

OpenStudy (anonymous):

i did 70- 30% is 21 is that right ?

OpenStudy (anonymous):

30%*70 \[\frac{ 30 }{ 100 }*70= 21\]

OpenStudy (anonymous):

how did u do that ? can u show me

OpenStudy (anonymous):

yo tutor where did u goo???????????????

OpenStudy (anonymous):

lol shush u cant complain about free stuff :P any percentage value is actually the number divided by a 100 so lets say u have X% that will be x/100 30% is actually 30/100 = 0.3 you need 30% of the 70 you just multiply the 0.3 * 70 or 30/100*70 = 2100/100 = 21

OpenStudy (anonymous):

ohhh... so u dont want me to complain lolz.. hehehehe whos complaining i am learningg

OpenStudy (anonymous):

ok i got it thannnkk =P

OpenStudy (anonymous):

lol women complain both ways

OpenStudy (anonymous):

oh godddd.. true but i am not a complainer

OpenStudy (anonymous):

ok what is %40 of 50

OpenStudy (anonymous):

ahhh =(

OpenStudy (anonymous):

20

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes how?

OpenStudy (anonymous):

u really gonna work my retriceof aren't you?

OpenStudy (anonymous):

40/100 = 0.4 = 0.4 X 50 = 20

OpenStudy (anonymous):

;)